EVA Report #2 – April 17th

Crew 245 EVA Report 17-04-2022
SOL: 7
EVA #: 8
Author: Dr. Nadia Maarouf

Purpose of EVA: Samples collection for the geology, sample collection tool testing, and Helmet testing

Start time: 07:35
End time: 10:00

Narrative: Crew drove to Moon Overlook (zone D from the new map), the "Moon" from the old map, then (possibly) up to north Copernicus to zone A from the new map. we tested the new helmet and aborted the EVA due to a backpack malfunction (#3). We drove back to the hab fast but safe.

Destination: Moon Overlook
Coordinates (UTM NAD27 CONUS): 4254400 N, 516000 E
4256400 N, 517000 E
Participants: Nadia (Commander) Paolo (HSO). Luca and Benjamin
Road(s) and routes per MDRS Map: North on Cow Dung Road 0110 then Brahe Hwy1572 then Copernicus Hwy 0157
Mode of travel: Rover and Walking

Copyright © The Mars Society. All rights reserved. | Main Site